OpenStudy (anonymous):

Which of the following expressions is this one equivalent to? (x^4 - 2x^3 + x - 2) / (x^3 + 1) a. x^2 - x + 1 b. x-2 c. x^2 + 3x + 2 d. x^2 + 1

OpenStudy (anonymous):

You can factor out common terms \[\huge \frac{ x^4-2x^3+x-2 }{ x^3+1 } \implies \frac{ (x-2)(x^3+1) }{ (x^3+1) }\]

OpenStudy (anonymous):

\[\frac{ x^4-2x^3+x-2 }{ x^3+1 } \implies \frac{ (x-2)(x^3+1) }{ (x^3+1) }\]

OpenStudy (anonymous):

Oh I never thought of it that way! That is so much easier than what I was initially taught! So the answers is going to be x-2! Thank you so much for your help!!

OpenStudy (anonymous):

Np :)